OpenStudy (anonymous):

y varies directly with x. When y = 20, x = 50. Find x when y = 36.

OpenStudy (ivettef365):

you need to use formula y = kx, find k, substitute y with 20 and x with 50

OpenStudy (ivettef365):

y=kx 20=k(50) 20/50=k 2/5=k k=2/5 y=2x/5 36=2x/5 5/2 * 36=x 5* 18=x 90=x